what is i-100? i to the -100 power?
Answer by Alan3354(69443)

what is i-100?
i-100 is just that, or -100 + i
------------------
i to the -100 power?
= 1/(i^100) = 1/1
= 1
---------
i is the sqrt(-1), so i^4 = 1.
100 = 4*25, so i^100 = (i^4)^25 = 1^25 = 1